I NEED TO SET UP AN EQUATION FOR THIS WORD PROBLEM:
THE SUM OF TWO NUMBERS IS 18. THREE TIMES THE GREATER NUMBER EXCEEDS FOUR TIMES THE SMALLER NUMBER BY 5. FIND THE NUMBERS.

Answer by venugopalramana(3286)

let the smaller number be x and greater number be y.
sum of numbers = x+y = 18.......(1)
three times the greater number = 3*y
four times the smaller number = 4*x
three times the greater number exceeds four times the smaller number by 5
hence 3*y-4*x=5
-4*x+3*y=5 .........(2).we have 2 unknowns.let us eliminate one unknown say x.for this multiply eqn.(1) with 4 and add to eqn.(2).
eqn.(1)*4 gives us ....4*x+4*y=4*18=72.......(3)adding to eqn.(2)
-4*x+3*y=5......(2) we get
4*y+3*y=7*y=72+5=77
y=77/7=11 substituting in 1 we get
x+11=18 or x=18-11=7